o, Nhat is-the Sum of he Finite arthmche Seriks ST
df 1. What is the simplest form of the radical expression? 625x20y8 2. How can you write the expression with a rationalized denominator? V49 2 3. What is the solution to the equation? 16 = (x-2 ) Page 14 a +
4Let f(x)-x+7 and gx)Fd((-3) S. Graph the function y 2Vx +4 10 8 6 10 10 6 -10 6. Is the relationship between the variables in this table a direct variation, an inverse variation, or neither. If it is a direct or inverse v write a function to model it. Page 2 4Q +
.pdf 7. A group of students are going to attend a concert. The cost per person varies inversely and depends on the number of people attending the concer. It will cost $75 per ticket if 2000 people attend. How much will the tickets be if 5000 people attend? 8. Wite the equation for the translation of y that has asymptotes at z -3 nd y 17 9. Describes the discontinuities and asymptotes for the rational function. y (x-3)(Cx+2) x+4)x-4)(x+2) 10. What is the quotient in simplified form? State any restrictions on the variable x2 -36 (6- x) x2 +3x-28 x +5x-36 Page 3 14+ FIO F12
odf 11. What is the 70th term of the sequence that begins -3, 1,5,9... 12. The 3d term in a geometric sequence is 40. The S term is 10. What the possible values of the 4h term? 13. Find the sum of the following series (4n - 9) Page 44
o.pdf 14. What is sy for 2+4+8+16+.. 15. Tyler begins funding a college education for his son. He saves $1000 the first year. Each year he saves 5% more than the previous year. How much money will be in the college fund at the end of 18 years? Page 4Q+
mPsaU8iMauKGL02wwMO/view rite a paragraph about what you arned and/or questions about what Period: ou didnt understand Date: oe Sompele iworke Unit b p215 Decades of College Dre Questions/Main Ideas: Notes: Show uovr work here Calodlate each erobabilitu, aiven that Page 2 5Q+
Answer #1

Solution:

(10). The series in not an arithmetic series. There is some misprint.

(1). 625120 у

=sqrt{25 imes 25 imes x^{10} imes x^{10} imes y^4 imes y^4}

=sqrt{(25)^2 imes (x^{10})^2 imes (y^4)^2}

(25) × (2.10) × (y4)

25104

(2).  

7 V49

rac{sqrt[3]{7}}{sqrt[3]{7^2}}

Multiplying by sqrt[3]{7} on nominator as well as denominator.

=rac{sqrt[3]{7} imes sqrt[3]{7}}{sqrt[3]{7^2} imes sqrt[3]{7}}

7

=rac{sqrt[3]{49}}{sqrt[3]{7^3}}

=rac{sqrt[3]{49}}{7}

This is the simplified form.

(3).

16=(x-2)^{rac{2}{3}}

Powering 3/2 on both sides

16^{rac{3}{2}}=[(x-2)^{rac{2}{3}}]^{rac{3}{2}}

(4^2)^{rac{3}{2}}=[(x-2)^{rac{2}{3}}]^{rac{3}{2}}

using the property, (x^{a})^{b}=x^{a imes b} . SO,

4^{2 imes rac{3}{2}}=(x-2)^{(rac{2}{3} imes rac{3}{2})}

4^3=x-2

64=x-2

x=64+2

x=66

(4).  

f(x)=x^2+7

g(x)=rac{x-3}{x}

(fog)(x)=f(rac{x-3}{x})

(fog) (x) = (-)2 + 7

(fog)(x)=rac{(x-3)^2}{x^2}+7

(fog)(x)=rac{x^2+9-6x}{x^2}+7

(fog)(-3)=rac{(-3)^2+9-6(-3)}{(-3)^2}+7

(fog)(-3)=rac{9+9+18}{9}+7

(fog)(-3)=rac{36}{9}+7

(fog)(-3)=4+7

(fog)(-3)=11